OpenStudy (jjuden):

how do i find the square root of 724 i totally forgot and studying for final

OpenStudy (misty1212):

Hi!!

OpenStudy (jjuden):

hello

OpenStudy (misty1212):

you mean without a calculator right? write in simplest radical form?

OpenStudy (jjuden):

yes

OpenStudy (jjuden):

yes

OpenStudy (misty1212):

724 is divisible by 4

OpenStudy (misty1212):

in fact it is \(724=4\times 181\) so \[\sqrt{724}=\sqrt{4\times 181}=\sqrt4\sqrt{181}=2\sqrt{181}\]